### v2

@@ 4 Algorithm and its Properties
The time complexity grows sharply with $p$ when using enumeration to solve SLIDE. Therefore, to enhance the practical usability of SLIDE, we have presented an iterative algorithm in Section 4.1. The subsequent Section 4.2 will exhibit both the computational and statistical properties of this algorithm.
@@ 4.1 Algorithmic Description
The inner loop employs the splicing technique proposed by Zhu et al. (2020). The idea behind the splicing technique is simple to grasp. For a given node $i$, if we have a current best guess for its neighbors’ topology, we can improve the quality of our estimate by including the nodes that are important to improve ${\mathcal{L}}_{n}({J}_{i})$ and eliminating the ones that have less importance. Below, we provide a formal characterization of the importance. Suppose the neighborhood $\mathcal{A}\subseteq V\setminus\{i\}$ and the interaction $\widehat{{J}_{i}}=\operatorname*{arg\,max}\limits_{{\mathrm{supp}}({J}_{i})={\mathcal{A}}}{\mathcal{L}}_{n}({J}_{i})$ are our current guess, we introduce two types of importance: (i) the backward importance $\widehat{{J}}_{ij}^{2}$ and (ii) the forward importance $\widehat{{\eta}}_{j}^{2}$ where $\widehat{{\eta}}_{j}=\nabla_{j}{\mathcal{L}}_{n}(\widehat{{J}}_{i})$.
@@
$\begin{split}\mathcal{S}_{s,1}&=\{j\in\mathcal{A}:\widehat{{J}}^{2}_{ij}\text{is one of the last $s$ backward importance}\},\\ \mathcal{S}_{s,2}&=\{j\in\mathcal{A}^{c}:\widehat{{\eta}}^{2}_{j}\text{ is oneof the top $s$ forward importance}\}.\end{split}$ (5)
@@
The main idea of middle loop is to determine the optimal neighbor size $d$ for the fixed node $i\in V$ that guarantees a good balance between the system’s accuracy and its complexity. In the literature, the balance is generally measured by information criterion such as Akaike information criterion and Bayesian information criterion (Akaike, 1974; Schwarz et al., 1978). In this paper, we propose to employ a generalized information criterion (GIC) [^2] Several information criteria are referred to as GIC in the literature. We defer the discussion with them in Supplementary Materials. . Specifically, given the interaction of the $i$-th node returned by the inner loop under the neighbor size $d_{i}$, denoted as $\widehat{{J}}_{i,d_{i}}$, GIC assesses $\widehat{{J}}_{i,d_{i}}$ by:
$\textup{GIC}(\widehat{{J}}_{i,d_{i}})\coloneqq{\mathcal{L}}_{n}(\widehat{{J}}_{i,d_{i}})-d_{i}\log p\log\log n,$
where the term $d_{i}\log p\log\log n$ encourages using fewer neighbors to explain the data so as to prevent overfitting caused by the incorrect inclusion of neighbors that weakly interact with node $i$. The proposed GIC is carefully designed to ensure consistent structure recovery. Specifically, $\log(p)$ is introduced to address overfitting in high-dimensional settings where $p$ is large. Besides, it is important to note that our proof requires an additional diverging sequence beyond $\log(p)$. To ensure the sequence diverges slowly, we set it to $\log\log(n)$. To get the optimal number of neighbors, the middle loop selects a $d_{i}$ from $\{0,1,\ldots,d_{\max}\}$ that maximizes $\textup{GIC}(\widehat{{J}}_{i,d_{i}})$.
@@
Two features of Algorithm 1 make it even more practically interesting. First, the algorithm exploits the fact that the number of unknown best neighbors from which to choose is discrete. Selecting a discrete neighbor’s number is more interpretable than choosing a continuous tuning parameter as the approaches in the literature. It is even more useful when there is a priori knowledge or expertise about the actual number of neighbors. For LASSO-type methods, this information has to be leveraged in a post-hoc manner.
Furthermore, during iterations, the proposed algorithm estimates the interactions between nodes without shrinking them, leading to an unbiased estimate of the interactions. Consequently, our algorithm gives more accurate estimation of the interactions, especially in the small-sample regime, which can be witnessed by the smallest interaction estimation errors in Figure 3.
Finally, in most cases, it is also challenging to know a priori what the smallest interaction ${\lambda}$ is. Notably, this procedure implicitly determines $\lambda$ in a data-driven way and gives a proper estimation for model complexity.
@@
5:         $\triangleright$ Inner loop: iteratively employing the splicing
@@
20:     $\widehat{{J}}_{ij}\leftarrow\frac{1}{2}(\widehat{{J}}_{ij}+\widehat{{J}}_{ji})\mathbb{I}(\frac{1}{2}|\widehat{{J}}_{ij}+\widehat{{J}}_{ji}|\geq\tau)$triangleright $\tau={\lambda}/2$ if ${\lambda}$ is known; otherwise $\tau=0$.
@@
  • Initialization strategy. In Step 4, the initialization of the inner loop sets the initial guess of $d$ neighbors as the union of (i) the previously estimated $d-1$ neighbors and (ii) the node with the highest forward importance. Notice that, the initial neighbor set $\mathcal{A}^{(0)}$ includes the previous estimated $d-1$ neighbors, i.e., ${\mathrm{supp}}(\widehat{J}_{i,d-1})$. This is a so-called warm-start initialization, a technique frequently employed in penalized regression to enhance empirical performance (see, e.g., Friedman et al., 2010; Zhu et al., 2022). Further, using $\{\operatorname*{arg\,max}\limits_{i}({\eta}_{i,d-1})^{2}\}$ bears a similar idea of marginal variable screening methods (Fan and Lv, 2008). To see this, consider the scenario where $d=1$ and $\mathcal{A}^{(0)}={\mathrm{supp}}(\widehat{J}_{i,0})=\emptyset$. In this scenario, we have ${\eta}_{i,0}=\left(\sum\limits_{r=1}^{n}{}^{(r)}_{1}{}^{(r)}_{i},\ldots,\sum\limits_{r=1}^{n}{}^{(r)}_{i-1}{}^{(r)}_{i},\sum\limits_{r=1}^{n}{}^{(r)}_{i+1}{}^{(r)}_{i},\ldots,\sum\limits_{r=1}^{n}{}^{(r)}_{p}{}^{(r)}_{i}\right),$ which records the empirical covariance between the $j$-th node and the $i$-th node (where $i\neq j$). Therefore, $\mathcal{A}^{(0)}$ is the node with the largest magnitude of covariance with the $i$-th node.
@@ 4.2 Theoretical Guarantees on the Algorithm
    $\Lambda_{i,d}({\bf Z})\coloneqq\Big{\{}\widehat{{J}}_{i}({\bf Z}):\,\widehat{{J}}_{i}({\bf Z})=\operatorname*{arg\,max}_{{\mathrm{supp}}({J}_{i})=\mathcal{A}}{\mathcal{L}}_{n}({J}_{i}),\forall\mathcal{A}\text{ s.t. }|\mathcal{A}|\leq d\Big{\}}\bigcup\Big{\{}{J}^{*}_{i}\Big{\}},$
@@
    $P\left\{\max_{i\in V}\max_{{J}_{i}\in\Lambda_{i,d_{\max}}({\bf Z})}\|{J}_{i}\|_{1}\leq c{\gamma}\right\}<\varepsilon_{0}=o(1).$
@@
Condition (C1) places a restriction on the magnitude of SLIDE across all possible models, since the curvature of PL may vanish at some points far away from the origin. In the context of generalized linear models, similar conditions with the same purpose are commonly encountered and are intended to avoid infinitely large or small variance of $y_{i}$, as well as ensure the existence of the Fisher information for statistical inference (Rigollet, 2012; Fan and Tang, 2013).
For arbitrary support size $d$, we define ${B}_{d}\coloneqq\max\limits_{\mathcal{A}:|\mathcal{A}|\leq d}\lambda_{\max}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{A}}^{\top}])$, ${b}_{d}\coloneqq\min\limits_{\mathcal{A}:|\mathcal{A}|\leq d}\lambda_{\min}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{A}}^{\top}])$, and ${\mu}_{d}\coloneqq\max\limits_{\mathcal{A},\mathcal{B}:\begin{subarray}{c}|\mathcal{A}|\leq d,\,|\mathcal{B}|\leq d,\mathcal{A}\cap\mathcal{B}=\varnothing\end{subarray}}\sigma_{\max}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{B}}^{\top}])$.
@@
    Denote $\gamma_{d}\coloneqq\frac{C_{1}(d,{\gamma},\Delta)}{(1-\Delta)C_{2}(d,{\gamma},\Delta)}$ where
    $C_{1}(d,{\gamma},\Delta)$ $=\frac{4(1+\Delta)^{7}\mu_{d}^{2}}{(1-\Delta)^{6}e^{-4c{\gamma}}b_{d}^{2}}\left(\frac{B_{d}}{2}+\frac{\mu_{d}^{2}}{e^{-2c{\gamma}}b_{d}}+\frac{B_{d}\mu_{d}^{2}}{2e^{-4c{\gamma}}b_{d}^{2}}\right)\left(2+\frac{2\mu_{d}}{e^{-2c{\gamma}}b_{d}}+\frac{\mu_{d}^{2}}{e^{-4c{\gamma}}b_{d}^{2}}\right),$
    $C_{2}(d,{\gamma},\Delta)$ $=\frac{(1-\Delta)}{2}{e^{-2c{\gamma}}b_{d}}-\frac{(1+\Delta)^{3}}{(1-\Delta)^{2}}\left(\frac{\mu_{d}^{2}}{e^{-2c{\gamma}}b_{d}}+\frac{B_{d}\mu_{d}^{2}}{2e^{-4c{\gamma}}b_{d}^{2}}\right),$
    Suppose that $\exists\Delta\in(0,1)$, such that $10\leq\gamma_{d_{\max}}<$ 1.
@@
Condition (C2) is a technical requirement. The quantities ${B}_{d}$, ${b}_{d}$ and ${\mu}_{d}$ are widely used in modeling high-dimensional data. They restrict the correlation among a small number of variables and, in turn, guarantee the identifiability of the true active set. ${B}_{d}$ and $e^{-2c{\gamma}}{b}_{d}$ are actually upper and lower bounds for the spectrum of the Hessian matrix over supports with a restricted cardinality. Condition (C2) is satisfied when they are closed enough and ${\mu}_{d}$ is sufficiently small. Specifically, in an asymptotic point of view, in order to guarantee that $C_{2}(d,{\gamma},\Delta)>0$, ${\mu}_{d}$ is required to decay exponentially with ${\gamma}$. Intuitively, this requires PL to be both smooth and convex enough, and column vectors of the design matrix are nearly orthogonal. In particular, this condition holds trivially if the design matrix has orthogonal columns.
@@
We now present the statistical guarantee of Algorithm 1. The following theorem asserts that the solution of the middle loop consistently identifies the underlying neighbors with high probability.
Theorem 2 (Consistency of Nodewise Neighbor Selection)
Assume Conditions (C1)-(C3) hold and the support size $d^{*}\coloneqq|{\mathrm{supp}}({J}_{i}^{*})|\leq d_{\max}$. Denote $\widehat{{J}}_{i,\widehat{d}}$ as the $j$-th column of matrix $\widehat{{J}}$ outputted by Algorithm 1 and $\widehat{\mathcal{A}}_{\widehat{d}}={\mathrm{supp}}(\widehat{{J}}_{i,\widehat{d}})$. Then there exists a universal constant $K>0$ such that, if
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}n>\max\Bigg{\{}\exp\left\{\exp\left[(1+\Delta)Ke^{4{\gamma}}(1+\frac{\delta}{d^{*}})\right]\right\},$ (6)
$\frac{\kappa d_{\max}\log p}{(1-\gamma_{d_{\max}})(1-\Delta)C_{2}(d_{\max},{\gamma},\Delta){\lambda}^{2}}\log\left[\frac{\kappa d_{\max}\log p}{(1-\gamma_{d_{\max}})(1-\Delta)C_{2}(d_{\max},{\gamma},\Delta){\lambda}^{2}}\right]\Bigg{\}},$
then the true active set is selected with high probability:
$P(\widehat{\mathcal{A}}_{i,\widehat{d}}={\mathrm{supp}}({J}_{i}^{*}))\geq 1-\varepsilon_{0}-\varepsilon_{1}(d_{\max})-\varepsilon_{2}(d_{\max})-\varepsilon_{3}(d_{\max})-\varepsilon^{\prime}_{4}(d_{\max})-d_{\max}p^{-\delta}.$
Here $\varepsilon_{1}(d)=2p^{2}\exp\Big{\{}-\frac{n\Delta^{2}{B}_{d}^{2}}{2d^{2}}\Big{\}}$, $\varepsilon_{2}(d)=2p^{2}\exp\Big{\{}-\frac{n\Delta^{2}{\mu}_{d}^{2}}{2d^{2}}\Big{\}}$, $\varepsilon_{3}(d)=2p^{2}\exp\Big{\{}-\frac{n\Delta^{2}{b}_{d}^{2}}{2d^{2}}\Big{\}}$ and $\varepsilon_{4}(d)=2p\exp\Big{\{}-\frac{n{\lambda}^{2}t^{2}(d,{\gamma},\Delta)}{8d}\Big{\}}$, $\delta>0$ is an arbitrary constant, and $\varepsilon_{4}^{\prime}(d)=2p\exp\Big{\{}-\frac{nt^{\prime}(d,{\gamma},\Delta,{\lambda},(\log{p}/n)^{\frac{1}{3}})^{2}}{8d}\Big{\}}$ where $t(d,{\gamma},\Delta)$ and $t^{\prime}(d,{\gamma},\Delta,{\lambda},(\log{p}/n)^{\frac{1}{3}})$ are constants whose explicit forms are provided in Supplementary Materials.
Since the neighbor sets is consistently detected, we can easily establish the property of parameter estimation.
@@
Under the same conditions and notations in Theorem 2, the $\ell_{2}$ error bound of $\widehat{{J}}$ returned by Algorithm 1 satisfies:
$\|\widehat{{J}}_{i}-{J}^{*}_{i}\|_{2}\leq\frac{t^{\prime}(d,{\gamma},\Delta,{\lambda},(\log{p}/n)^{\frac{1}{3}})}{(1-\Delta)e^{-2c{\gamma}}b_{d}},\;\textup{for}\;i=1,\ldots,p.$
Applying the union bound, it is straightforward to obtain the following corollary, which ensures that Algorithm 1 consistently reconstructs the whole graph.
Corollary 2 (Consistency of Structural Recovery for Ising models)
Under the same conditions and notations in Theorem 2, then with probability at least $1-\varepsilon_{0}-\varepsilon_{1}(d_{\max})-\varepsilon_{2}(d_{\max})-\varepsilon_{3}(d_{\max})-p\varepsilon^{\prime}_{4}(d_{\max})-d_{\max}p^{1-\delta}$, Algorithm 1 will select the true active set, i.e.,
${\mathrm{supp}}(\widehat{{J}})={\mathrm{supp}}({J}^{*}).$
The next theorem claims that, in a high probability sense, Algorithm 1 has a polynomial time complexity.
@@
Assume Conditions (C1)-(C3) hold and Algorithm 1 successfully recover the true graphical structure ${\mathrm{supp}}(\widehat{{J}})={\mathrm{supp}}({J}^{*})$ (by Corollary 2, this event is true with a large probability for sufficiently large sample size $n$). Then the computational complexity of Algorithm 1 for a given $d_{\max}\geq\max\limits_{i\in V}|{\mathrm{supp}}({J}^{*}_{i})|$ is
$O\Big{(}\big{(}\frac{d_{\max}}{1-\gamma_{d_{\max}}}\ln\frac{n{\gamma}}{\min\limits_{i\in V}|{\mathrm{supp}}({J}^{*}_{i})|\log p\log\log n}+\frac{n{\gamma}}{\log p\log\log n}\big{)}\big{(}s_{\max}pnd_{\max}+s_{\max}p^{2}\big{)}\Big{)}.$
We claim that the computational complexity of Algorithm 1 scale as a polynomial in $(n,p,d)$. In other words, under certain conditions, our algorithm solves the intractable problem of learning sparse Ising model in a polynomial time. To see this, we only need to verify that $({B}_{d},{\mu}_{d},{b}_{d})$ can be selected such that $\frac{1}{1-\gamma_{d}}=\frac{(1-\Delta)C_{2}}{(1-\Delta)C_{2}-C_{1}}$ scales at most polynomially with $(n,p,d)$. This is true since $\lambda_{\max}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{A}}^{\top}])\leq d$, $\sigma_{\max}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{B}}^{\top}])\leq d$, and $\lambda_{\min}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{A}}^{\top}])\geq\frac{e^{-2{\gamma}}}{d+1}$ where the last inequality is also presented in Lemma 7 of Vuffray et al. (2016). Note that the complexity grows exponentially fast as ${\gamma}$ increases, which is aligned with our expectation since the complexity is derived for the case that our algorithm successfully recovers the graphical structure, which, in turn, requires a sample size $n$ to scale exponentially with ${\gamma}$.

### v3

@@ 4 Algorithm and its Properties
The time complexity grows exponentially with $p$ when using exhaustive enumeration to solve SLIDE. Therefore, to enhance the practical usability of SLIDE, we present an iterative algorithm in Section 4.1. The subsequent Section 4.2 will exhibit both the computational and statistical properties of this algorithm.
@@ 4.1 Algorithmic Description
The inner loop employs the splicing technique proposed by Zhu et al. (2020). The idea behind the splicing technique is simple to grasp. For a given node $i$, if we have a current guess for its neighbors, we may improve the quality of our guess by including the nodes that are important to improve ${\mathcal{L}}_{n}({J}_{i})$ and eliminating the ones that have less important on ${\mathcal{L}}_{n}({J}_{i})$. Below, we provide a formal characterization of the importance. Suppose the neighborhood $\mathcal{A}\subseteq V\setminus\{i\}$ and the interaction $\widehat{{J}_{i}}=\operatorname*{arg\,max}\limits_{{\mathrm{supp}}({J}_{i})={\mathcal{A}}}{\mathcal{L}}_{n}({J}_{i})$ are our current guess, we introduce two types of importance: (i) the backward importance $\widehat{{J}}_{ij}^{2}$ and (ii) the forward importance $\widehat{{\eta}}_{j}^{2}$ where $\widehat{{\eta}}_{j}=\nabla_{j}{\mathcal{L}}_{n}(\widehat{{J}}_{i})$.
@@
[…]
@@
The main idea of middle loop is to determine the optimal neighbor size $d$ for the fixed node $i\in V$ that guarantees a good balance between the system’s accuracy and its complexity. In the literature, the balance is generally measured by information criterion such as Akaike information criterion and Bayesian information criterion (Akaike, 1974; Schwarz et al., 1978). In this paper, we propose to employ a generalized information criterion (GIC) [^2] Several information criteria are referred to as GIC in the literature. We defer the discussion with them in the Supplementary Materials. . Specifically, given the interaction of the $i$-th node returned by the inner loop under the neighbor size $d_{i}$, denoted as $\widehat{{J}}_{i,d_{i}}$, GIC assesses $\widehat{{J}}_{i,d_{i}}$ by:
$\textup{GIC}(\widehat{{J}}_{i,d_{i}})\coloneqq{\mathcal{L}}_{n}(\widehat{{J}}_{i,d_{i}})-d_{i}c_{n}\log p,$
where the term $d_{i}c_{n}\log p$ encourages using fewer neighbors to explain the data so as to prevent overfitting caused by the incorrect inclusion of neighbors that weakly interact with node $i$. The proposed GIC is carefully designed to ensure consistent structure recovery. Specifically, $\log p$ is introduced to address overfitting in high-dimensional settings where $p$ is large. In our algorithm, we set $c_{n}=\log\log n$ to ensure the sequence diverges slowly. To get the optimal number of neighbors, the middle loop selects a $d_{i}$ from $\{0,1,\ldots,d_{\max}\}$ that maximizes $\textup{GIC}(\widehat{{J}}_{i,d_{i}})$.
@@
Two features of Algorithm 1 make it even more practically intriguing. First, the algorithm exploits the fact that the number of unknown best neighbors from which to choose is discrete. Selecting a discrete neighbor’s number is more interpretable than choosing a continuous tuning parameter as the approaches in the literature. It is even more useful when there is a priori knowledge or expertise about the actual number of neighbors. For LASSO-type methods, this information has to be leveraged in a post-hoc manner. Furthermore, during iterations, the proposed algorithm estimates the interactions between nodes without shrinking them, leading to an unbiased estimate of the interactions. Consequently, our algorithm gives more accurate estimation of the interactions, especially in the small-sample regime, which can be witnessed by the smallest interaction estimation errors in Figure 3. Finally, in most cases, it is also challenging to know a priori what the smallest interaction $\lambda$} is. Notably, this procedure implicitly determines $\lambda$ in a data-driven way and gives a proper estimation for model complexity.
@@
[…]
@@
20:  $\widehat{{J}}_{ij}\leftarrow\frac{1}{2}(\widehat{{J}}_{ij}+\widehat{{J}}_{ji})\mathbb{I}(\frac{1}{2}|\widehat{{J}}_{ij}+\widehat{{J}}_{ji}|\geq\tau)$ $\triangleright$ $\tau={\lambda}/2$ if $\lambda$} is known; otherwise $\tau=0$.
@@
  • Initialization strategy. In Step 4, the initialization of the inner loop sets the initial guess of $d$ neighbors as the union of (i) the previously estimated d1d-1 neighbors and (ii) the node with the highest forward importance. Notice that, the initial neighbor set $\mathcal{A}^{(0)}$ includes the previous estimated d1d-1 neighbors, i.e., ${\mathrm{supp}}(\widehat{J}_{i,d-1})$. This is a so-called warm-start initialization, a technique frequently employed in penalized regression to enhance empirical performance (see, e.g., Friedman et al., 2010). Furthermore, using $\{\operatorname*{arg\,max}\limits_{i}({\eta}_{i,d-1})^{2}\}$ bears a similar idea of marginal variable screening methods (Fan & Lv, 2008). To see this, consider the scenario where $d=1$ and $\mathcal{A}^{(0)}={\mathrm{supp}}(\widehat{J}_{i,0})=\emptyset$. In this scenario, we have ${\eta}_{i,0}=\left(\sum\limits_{r=1}^{n}\real{}^{(r)}_{1}\real{}^{(r)}_{i},\ldots,\sum\limits_{r=1}^{n}\real{}^{(r)}_{i-1}\real{}^{(r)}_{i},\sum\limits_{r=1}^{n}\real{}^{(r)}_{i+1}\real{}^{(r)}_{i},\ldots,\sum\limits_{r=1}^{n}\real{}^{(r)}_{p}\real{}^{(r)}_{i}\right),$ which records the empirical covariance between the $j$-th node and the $i$-th node (where $i\neq j$). Therefore, $\mathcal{A}^{(0)}$ is the node with the largest magnitude of covariance with the $i$-th node.
@@ 4.2 Theoretical Guarantees on the Algorithm
    $\Lambda_{i,d}\coloneqq\Big\{\widehat{{J}}_{i}:\,\widehat{{J}}_{i}=\operatorname*{arg\,max}_{{\mathrm{supp}}({J}_{i})=\mathcal{A}}{\mathcal{L}}_{n}({J}_{i}),\forall\mathcal{A}\text{ s.t. }|\mathcal{A}|\leq d\Big\}\bigcup\Big\{{J}^{*}_{i}\Big\},$
@@
    $P\left\{\max_{i\in V}\max_{{J}_{i}\in\Lambda_{i,d_{\max}}}\|{J}_{i}\|_{1}\leq c{\gamma}\right\}<\varepsilon_{0}=o(1).$
@@
Condition (C1) places a restriction on the magnitude of SLIDE across all possible models, since the curvature of PL may vanish at some points far away from the origin. In the context of generalized linear models, similar conditions with the same purpose are commonly encountered and are intended to avoid infinitely large or small variance of $y_{i}$, as well as ensure the existence of the Fisher information for statistical inference (Rigollet, 2012; Fan & Tang, 2013).
For arbitrary support size $d$, we define ${B}_{d}^{0}\coloneqq\max\limits_{\mathcal{A}:|\mathcal{A}|\leq d}\lambda_{\max}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{A}}^{\top}])$, ${b}_{d}^{0}\coloneqq\min\limits_{\mathcal{A}:|\mathcal{A}|\leq d}\lambda_{\min}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{A}}^{\top}])$, and ${\mu}_{d}^{0}\coloneqq\max\limits_{\mathcal{A},\mathcal{B}:\begin{subarray}{c}|\mathcal{A}|\leq d,\,|\mathcal{B}|\leq d,\mathcal{A}\cap\mathcal{B}=\varnothing\end{subarray}}\sigma_{\max}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{B}}^{\top}])$.
@@
    For some small constant $c_{0}>0$, denote ${B}_{d}={B}_{d}^{0}+c_{0}$, ${m}_{d}=e^{-2c{\gamma}}({b}_{d}^{0}-c_{0})$ and ${\mu}_{d}={\mu}_{d}^{0}+c_{0}$, with constant $c$ as defined in (C1). There exists some constant $\Delta\in(0,1)$, such that $0\leq\nu({d_{\max}},{\gamma})<1$. Here $\nu(d,{\gamma})\coloneqq\frac{C_{1}(d,{\gamma})}{(1-\Delta)C_{2}(d,{\gamma})}$ and
    $C_{1}(d,{\gamma})=\frac{2{\mu}_{d}^{2}}{{m}_{d}^{2}}\left(\frac{{B}_{d}}{2}+\frac{{\mu}_{d}^{2}}{{m}_{d}}+\frac{{B}_{d}{\mu}_{d}^{2}}{2{m}_{d}^{2}}\right)\left(3+\frac{2{\mu}_{d}}{{m}_{d}}\right)^{2},\;\;C_{2}(d,{\gamma})=\frac{{m}_{d}}{2}-\frac{{\mu}_{d}^{2}}{{m}_{d}}-\frac{{B}_{d}{\mu}_{d}^{2}}{2{m}_{d}^{2}}.$
@@
Condition (C2) is a technical requirement. The quantities ${B}_{d}^{0}$, ${b}_{d}^{0}$ and ${\mu}_{d}^{0}$ are widely used in modeling high-dimensional data. They restrict the correlation among a small number of variables and, in turn, guarantee the identifiability of the true active set. ${B}_{d}$ and ${m}_{d}$ are actually upper and lower bounds for the spectrum of the Hessian matrix over supports with a restricted cardinality. Condition (C2) is satisfied when they are closed enough and ${\mu}_{d}$ is sufficiently small. Specifically, in an asymptotic point of view, in order to guarantee that $C_{2}(d,{\gamma})>0$, ${\mu}_{d}$ is required to decay exponentially with $\gamma$}. Intuitively, this requires PL to be both smooth and convex enough, and column vectors of the data matrix are nearly orthogonal. In particular, this condition holds trivially if the data matrix has orthogonal columns.
@@
We now present the statistical guarantee of Algorithm 1. The following theorem asserts that the solution of Algorithm 1 consistently recovers the underlying Ising model with high probability.
Theorem 2 (Consistency of Algorithm 1 in Reconstructing Ising Models)
Denote $\widehat{{J}}$ as the output of Algorithm 1. Assume Conditions (C1)-(C3) hold and that $c_{n}\in[(1+\Delta)K_{0}e^{4{\gamma}}(1+\delta),\frac{C_{2}(d_{\max},{\gamma}){\lambda}^{2}n}{d_{\max}\log p}]$ for some constant $K_{0}>0$ and arbitrary constant $\delta>1$. Suppose for some universal constant $K_{1}>0$
$n>\frac{K_{1}d_{\max}\log p}{[1-\nu(d_{\max},{\gamma})](1-\Delta)C_{2}(d_{\max},{\gamma}){\lambda}^{2}}\log\left[\frac{d_{\max}\log p}{[1-\nu(d_{\max},{\gamma})](1-\Delta)C_{2}(d_{\max},{\gamma}){\lambda}^{2}}\right],$ (6)
then the support set of ${J}^{*}$ is selected with high probability:
$P\left({\mathrm{supp}}(\widehat{{J}})={\mathrm{supp}}({J}^{*})\right)\geq 1-\varepsilon_{0}-\varepsilon_{1}(d_{\max})-\varepsilon_{2}(d_{\max})-\varepsilon_{3}(d_{\max})-\varepsilon_{4}(d_{\max})-d_{\max}p^{1-\delta}.$
Here $\varepsilon_{1}(d)=2p^{3}\exp\Big\{-\frac{nc_{0}^{2}}{2d^{2}}\Big\}$, $\varepsilon_{2}(d)=2p^{2}\exp\Big\{-\frac{n{\lambda}^{2}t^{2}(d,{\gamma})}{8d}\Big\}$, $\varepsilon_{3}(d)=2p^{2}\exp\Big\{-\frac{ne^{-4{\gamma}}({b}_{d}^{0}-c_{0})^{2}}{72d^{2}{B}_{d}^{2}}\Big\}$ and $\varepsilon_{4}(d)=2p^{2}\exp\Big\{-\Big(\frac{n\tilde{t}^{2}(d,{\gamma})}{8^{3}d^{4}}\Big)^{\frac{1}{3}}\Big\}$. The explicit forms of $t(d,{\gamma})$ and $\tilde{t}(d,{\gamma})$ are provided in the Supplementary Materials.
Remark 8
Theorem 2 helps reveal the sample complexity of Algorithm 1 (denoted as $n^{\textup{alg}}$) with respect to $\lambda$}, $p$, $d$ and $\gamma$}. It is easy to see that $n^{\textup{alg}}\;\propto\;{\lambda}^{-2}\log{\lambda}^{-2}$ and $n^{\textup{alg}}\;\propto\;\log p\log\log p$. And thus, for $\lambda$} and $p$, $n^{\textup{alg}}$ lies in the globally optimal region up to some logarithmic terms. The relationship between $n^{\textup{alg}}$ and $d$ is not directly evident, as it depends on two quantities $C_{2}(d,{\gamma})$ and $\nu(d,{\gamma})$ [^3] Here we set $d_{\max}=Cd$, where $C>1$ is a constant to simplify notations and make the results of Theorem 2 being comparable to Theorem 1.. By lower-bounding $C_{2}(d,{\gamma})$, upper-bounding $\nu(d,{\gamma})$, and assuming $\lambda$} is known (as the baselines in Table 1), we have $n^{\textup{alg}}\;\propto\;d^{6}$. A similar argument for $\gamma$} yields $n^{\textup{alg}}\;\propto\;e^{8(c+1){\gamma}}$ for some constant $c\geq 1$ defined in Condition (C1). Although there exists a gap between $n^{\textup{alg}}$ and $n^{*}$ in terms of $\gamma$} and $d$, this gap appears to be an artifact of the worst-case analysis. Indeed, Figures 12 show that the empirical sample complexity $n^{\textup{emp}}$ of Algorithm 1 matches—or even outperforms—the theoretical SLIDE estimator. For example, $n^{\textup{emp}}\;\propto\;e^{3.5{\gamma}}$ grows more slowly than SLIDE’s rate $e^{4{\gamma}}$.
Since the support set is consistently detected, we can easily establish the property of parameter estimation.
@@
Under the same conditions and notations in Theorem 2, with additional probability of error $\varepsilon$, the (squared) $\ell_{2}$ error bound of the $i$-th column of $\widehat{{J}}$ returned by Algorithm 1 satisfies:
$\|\widehat{{J}}_{i}-{J}^{*}_{i}\|^{2}_{2}\leq\frac{1}{{m}^{2}_{d}}\frac{16d}{n}\log\left(\frac{2p}{\varepsilon}\right)$
for all $i=1,\ldots,p$.
The next theorem claims that, in a high probability sense, Algorithm 1 has a polynomial-time complexity.
@@
Assume Conditions (C1)-(C3) hold and Algorithm 1 successfully recovers the true graphical structure ${\mathrm{supp}}(\widehat{{J}})={\mathrm{supp}}({J}^{*})$ (by Theorem 2, this event occurs with high probability for a sufficiently large sample size $n$). Then the computational complexity of Algorithm 1 for a given $d_{\max}\geq\max\limits_{i\in V}|{\mathrm{supp}}({J}^{*}_{i})|$ is
$O\Big(\big(\frac{d_{\max}}{1-\nu(d_{\max},{\gamma})}\ln\frac{n{\gamma}}{\min\limits_{i\in V}|{\mathrm{supp}}({J}^{*}_{i})|\log p\log\log n}+\frac{n{\gamma}}{\log p\log\log n}\big)\big(pnd^{2}_{\max}+d_{\max}p^{2}\big)\Big).$
We claim that the computational complexity of Algorithm 1 scales as a polynomial in $(n,p,d)$. In other words, under certain conditions, our algorithm solves the intractable problem of learning sparse Ising model in a polynomial time. To see this, we only need to verify that $\frac{1}{1-\nu(d,{\gamma})}=\frac{(1-\Delta)C_{2}}{(1-\Delta)C_{2}-C_{1}}$ scales at most polynomially with $(n,p,d)$. This is true since $\lambda_{\max}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{A}}^{\top}])\leq d$, $\sigma_{\max}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{B}}^{\top}])\leq d$, and $\lambda_{\min}(\mathbb{E}[Z_{\mathcal{A}}Z_{\mathcal{A}}^{\top}])\geq\frac{e^{-2{\gamma}}}{d+1}$ where the last inequality is also presented in Lemma 7 of Vuffray et al. (2016). Note that the complexity grows exponentially fast as $\gamma$} increases, which is aligned with our expectation since the complexity is derived for the case that our algorithm successfully recovers the graphical structure, which, in turn, requires a sample size $n$ to scale exponentially with $\gamma$}.